6. Adult tickets for a play cost $17 and child tickets cost $8. If there were 24 people at a performance and the theater collected $363 from ticket sales, how many children attended the play?
A) 4 children
B) 5 children
C) 6 children
D) 19 children

7. Solve the following system of linear inequalities by graphing. 3x + 4y £ 12 x + 3y £ 6 x ³ 0 y ³ 0
)

Solve the system by addition or substitution. 3x + 6y = 0
x = Solve the system by graphing. 2x + y = 4 x + y = 3
I need some guidance here, can anyone assist me?



Answer by tutorcecilia(2144)

($17)(x)+($8)(24-x)=$363
17x+192-8x=363
17x-8x=363-192
9x=171
x=171/9
x=19 adults
Childrens tickets = (24-x)=24-19=5
.
Check by plugging (x=19) back into the original equation.
.
I don't understand the symbols in Question #7.